solve the system of equations using the addition method
2x-3y=4
3x-5y=4

Answer by checkley77(12844)

2x-3y=4 multiply by -3
3x-5y=4 multiply by 2 & add
-6x+9y=-12
6x-10y=8
--------------
-y=-4
y=-4/-1
y=4 ans.
2x-3*4=4
2x-12=4
2x=4+12
2x=16
x=16/2
x=8 ans.
Proof:
3*8-5*4=4
24-20=4
4=4